OpenStudy (sachi):

form a quadratic polynomial p(x) with 4 and -2 as the sum and product of its zeroes respectively

OpenStudy (sachi):

pls help me

OpenStudy (dls):

Sum of Roots = \(\Huge \frac{-b}{a}\) Product of Roots =\(\Huge \frac{c}{a}\) Sum = 4 and product = -2 (Given) Lastly, \[\Huge p(x)=ax^2+bx+c\] Put all the values and solve :)

OpenStudy (sachi):

thanku ummmmmm now can I ask 1 more ques

OpenStudy (dls):

After you close this one.